# From Cursed to Competitive: Closing the ZO–FO Gap via Input-to-State Stability

### 28 Apr 2026 v1

@@ Appendix D Complementary remarks on the results of Section 3
We note that the expected performance of ZO-GD for strongly convex objective functions in minimisation problems has been analysed in [28, Thm. 8]. Compared to Theorem 3, the result in [28, Thm. 8] exhibits an explicit dependence of the iteration complexity on the dimension, and does not account for the variance bound $\sigma$ or the number of sampled directions $t$. Moreover, the prescribed step size scales inversely with $n$, which is not consistent with standard GD. In contrast, Theorem 3 shows that ZO-GD, in expectation, achieves the same decay rate as GD and converges to a neighbourhood of the GD fixed point, whose size depends on $\sigma$ and can be manipulated via $h$, $\mu$, and $t$. Furthermore, for any step size consistent with GD, one can select $\mu$ and $t$ such that the perturbation is arbitrarily small and closely track GD.
@@
where $c=2h\beta L_{1}/(\beta+L_{1})\in(0,1)$ is the GD contraction constant derived in [27, Thm 2.1.15]. Using Young’s inequality, for any $\varepsilon>0$,
@@
If, in the proof of Lemma 5, one bounds $E[\|e_{k}\|]$ instead of $E[\|e_{k}\|^{2}]$, the resulting requirement becomes $h\in\bigl(0,\frac{1-c}{2L_{1}}\bigr]$, which implies $\lambda>L_{1}$ and hence guarantees that the regularised function is $L_{1}$-strongly convex; in that case, the results of Section 3.1 apply directly. The bound obtained in Lemma 5 is considerably more general, as it covers regimes where the regularised objective is not strongly convex.
Remark 10.
Following a similar argument to the proof of Lemma 5, one can show that for $h\in\bigl(0,\frac{\sqrt{2(1+c^{2})-2c}}{2L_{1}}\bigr)$, the same bound as in (36) holds but with $2L_{1}^{2}h^{2}$ in place of $8L_{1}^{2}h^{2}$. This tighter bound permits a nearly twice as large admissible range for $h$; however, the corresponding choice of $\lambda$ satisfies $\lambda>\frac{L_{1}}{2}$, which is precisely the threshold at which the regularised objective becomes strongly convex. Consequently, $\lambda$ cannot be made arbitrarily small, and the minimiser of the regularised problem cannot be brought arbitrarily close to that of the original problem. This reflects the well-known trade-off between performance and robustness.
@@ E.1 Additional experiments for the quadratic objective
Limitations of the number of sampled directions compensation.
To demonstrate that increasing $t$ alone cannot fully compensate for a large $\mu$, we consider three configurations with $\mu=5\times 10^{-1}$ fixed: (i) $h=10^{-5}$, $t=1$; (ii) $h=10^{-7}$, $t=1$; and (iii) $h=10^{-5}$, $t=20$. The results are shown in Figure 7. In case (i), ZO-GD diverges while GD converges. Reducing $h$ in case (ii) shrinks the perturbation and restores convergence. Increasing $t$ in case (iii) also reduces the perturbation, but since $t$ does not affect the smoothing bias term in Lemma 2, it cannot fully compensate for a large $\mu$.
Figure 7: Parameter analysis: limitations of the number of sampled directions compensation.

### 01 May 2026 v2

@@ Appendix D Complementary remarks on the results of Section 3
We note that the expected performance of ZO-GD for strongly convex objective functions in minimisation problems has been analysed in [27, Thm. 8]. Compared to Theorem 3, the result in [27, Thm. 8] exhibits an explicit dependence of the iteration complexity on the dimension, and does not account for the variance bound $\sigma$ or the number of sampled directions $t$. Moreover, the prescribed step size scales inversely with $n$, which is not consistent with standard GD. In contrast, Theorem 3 shows that ZO-GD, in expectation, achieves the same decay rate as GD and converges to a neighbourhood of the GD fixed point, whose size depends on $\sigma$ and can be manipulated via $h$, $\mu$, and $t$. Furthermore, for any step size consistent with GD, one can select $\mu$ and $t$ such that the perturbation is arbitrarily small and closely track GD.
@@
where $c=2h\beta L_{1}/(\beta+L_{1})\in(0,1)$ is the GD contraction constant derived in [28, Thm 2.1.15]. Using Young’s inequality, for any $\varepsilon>0$,
@@
We note that the step-size condition in Lemma 5, combined with the parametrisation $\lambda=\frac{1-c}{2h}$, implies $\lambda>\frac{L_{1}}{2}$, so the regularised objective is strongly convex across the entire admissible parameter range. Nevertheless, the $L_{2}$ regularisation framework remains valuable: it provides a systematic mechanism for applying the ISS analysis to objectives that are not originally strongly convex, with explicit control over the trade-off between the regularisation bias and the convergence neighbourhood radius through the parameters $c$, $h$, $\mu$, and $t$. Extending the analysis to parameter regimes where the regularised objective is not strongly convex is an interesting direction for future work.